# Di-NeRF: Distributed NeRF for Collaborative Learning with Unknown Relative Poses

### v1

@@ Abstract
Collaborative mapping of unknown environments can be done faster and more robustly than a single robot. However, a collaborative approach requires a distributed paradigm to be scalable and deal with communication issues. This work presents a fully distributed algorithm enabling a group of robots to collectively optimize the parameters of a Neural Radiance Field (NeRF). The algorithm involves the communication of each robot’s trained NeRF parameters over a mesh network, where each robot trains its NeRF and has access to its own visual data only. Additionally, the relative poses of all robots are jointly optimized alongside the model parameters, enabling mapping with unknown relative camera poses. We show that multi-robot systems can benefit from differentiable and robust 3D reconstruction optimized from multiple NeRFs. Experiments on real-world and synthetic data demonstrate the efficiency of the proposed algorithm. See the website of the project for videos of the experiments [^1] https://sites.google.com/view/di-nerf/home.
@@ I Introduction
There is an increasing demand for robots to collaborate on complex tasks, such as mapping an unknown environment. Centralized approaches for the coordination of the robots or processing data face scalability issues, vulnerability to central node failures, and the risk of communication blackouts. A true collaborative robotic system needs to work in a distributed manner [1]. Further, building a high-quality representation in collaborative mapping is needed to make an informed decision. The representations also need to be compact for easy sharing. Neural Radiance Field (NeRF) [2] enables such a representation for a single robot, leveraging advancements in neural networks. Extending the capabilities of NeRF to facilitate multi-robot NeRF in a distributed manner emerges as a natural progression. This extension allows for the collaborative development of high-quality representations of unknown environments without dependence on a central node.
Performing distributed NeRF requires addressing several problems inherent to multi-robot systems [3]. Firstly, determining the relative poses of robots, preferably without requiring them to rendezvous. This is needed to enable the fusions of the local maps (i.e. maps of individual robots) into a global map. The relative poses can be determined either via line-of-sight rendezvous [4], which necessitates motion coordination, or by identifying overlaps within local maps [5][6], requiring additional processing. In large-scale applications, avoiding line-of-sight rendezvous is preferred to prevent further planning constraints. The second challenge is deciding what information to share among robots. While sharing unprocessed raw visual data is resource-intensive and raises privacy concerns, sharing compact neural models is a more viable option. However, sharing maps introduces complexity to the subsequent challenge. The third challenge is merging representations, requiring the development of a method to unify local neural maps into a global map. Merging neural representations is more intricate than merging geometric representations. These challenges must be systematically addressed in a distributed manner to ensure scalability and robustness against node failure.
@@
There are many methods for distributed/centralized geometric multi-robot mapping and localization [7][4][8]. For learned methods, federated [9] and distributed learning [1] are common for overcoming the issues associated with centralized learning. In these methods, the training process is decentralized. Each robot performs its training, and the results are aggregated in a central node [9]. The need for a central node still constrains the scalability of such methods. A recent work using multiple agents for NeRF is Block-NeRF [10], composed of independent local maps, i.e. NeRF block. In Block-NeRF, the relative poses of the blocks are known to a degree and the task is done in a centralized manner. Another closely related work is DiNNO [11], where distributed mapping is performed using NeRF. DiNNO performs 2D mapping, and the assumption is that the relative poses of the robots are known.
In this paper, we present Di-NeRF (Distributed Neural Radiance Fields), an algorithm that builds on the Consensus Alternating Direction Method of Multipliers (C-ADMM) as a versatile distributed optimization method for multi-robot systems. In Di-NeRF, each robot starts building its NeRF by relying on its own data. By integrating NeRFs from other robots, each robot will build a global NeRF model with a quality comparable to a fully centralized method. The advantages of Di-NeRF include the robots do not know their initial poses, and there is no robot rendezvous. The robots do not send raw data over wireless networks, protecting privacy and optimizing wireless bandwidth usage, as illustrated graphically in Fig. 1. In Di-NeRF, robots alternate between local optimization of an objective function and communication of intermediate network weights over the wireless network. Di-NeRF can consider different communication graphs (e.g. fully connected, circular, and ring connectivity). The key contributions of Di-NeRF include i) developing fully distributed optimization for 3D reconstruction with RGB images as input, and a NeRF as the backend, enabling fusing NeRFs in training not the rendering process, and ii) optimization of the relative poses of the robots, eliminating the need for a global coordinate system for all the robots and removing the need to know the prior relative poses.
@@ II Literature Review
Multi-robot localization and mapping algorithms utilize various representations such as sparse landmarks [12, 13], dense geometric maps [14], object classes [15, 16], and semantics [17, 18, 19]. With the advent of neural radiance field [2] and its variants [20, 21, 22, 23, 24, 25], there are efforts to use neural maps for collaborative localization/mapping [10, 26, 27].
Emerging radiance fields, including non-neural representations such as Plenoxels [28], and neural representations, such as NeRF [2] have revolutionized localization and mapping algorithms. Nice SLAM [29], iMAP [30], NeRF-SALM[31], and [32] are using such representations. The early versions of NeRF representations [25, 23, 33, 34, 35, 36, 37] required pose information often generated via Structure-from-Motion (SfM) algorithms, e.g. COLMAP [38]. This limitation was Later addressed in [20, 24, 39, 40, 41, 42], by jointly optimizing camera poses and the scene. It was also shown that a camera view can be localized in a NeRF map, as shown in iNeRF [43].
Extending NeRF techniques to multi-robot scenarios has also been proposed in Block-NeRF [10], NeRFuser [44], and NeRFusion [45]. These methods use blending techniques to render an image, and often this is done in a centralized manner, performing inference on multiple NeRF models. Furthermore, it is assumed that there is prior knowledge about the relative pose of the robots. A similar research trend is federated learning [26, 27, 46], where the training is done in a decentralized manner, and each node learns a common NeRF in parallel. Then the weights are transferred to a server for aggregation. A closely related work is DiNNO [11], a distributed algorithm that enables multiple robots to collaboratively optimize the parameters of a neural network to learn a 2D map. Each robot maintains its version of a neural network model and has access only to its data. They communicate the models to achieve a consensus; however, prior knowledge of each robot’s trajectory is needed for the algorithm to update the models. The core of the algorithm is a distributed optimization algorithm known as C-ADMM [1, 47].
Figure 2: Three connected robots share their local NeRF to build a global NeRF. Each robot adjusts its weight compared to receiving weights using C-ADMM, and all robots except robot $R^{g}$ (Robot#1, the robot whose coordinate is considered as the global coordinate) optimize for the relative pose $T_{i}^{g}$.
@@ III Distributed Neural Radiance Field (Di-NeRF)
Here the problem is defined followed by the proposed method for relative pose estimation and distributed NeRF. The overall pipeline for Di-NeRF is shown in Fig. 2. Di-NeRF is a fully distributed algorithm for 3D reconstruction using RGB images. We assume the relative poses of the robots are unknown. With Di-NeRF, each robot has its local coordinate and only shares the learned models with other robots to achieve a consensus while optimizing for the relative poses.
@@ III-A Problem Statement
Given a set of images captured from a few viewpoints of a scene with the associated local camera parameters, the goal of Di-NeRF is to build a scene representation that enables all the robots to generate realistic images from novel viewpoints. Furthermore, since the poses of each robot are defined in a local coordinate system, it is necessary to perform relative pose optimization to build a consistent model that accurately represents the entire scene. The problem is expressed as:
$\theta,\textbf{T}=\underset{\theta,\textbf{T}}{\operatorname{argmin}}\sum_{i\in\mathcal{V}}{L_{i}(\theta,T^{g}_{i},I_{i},\pi_{i})},\leavevmode\nobreak\ \textbf{T}=\{T^{g}_{1},..,T^{g}_{N}\},$ (1)
where $L_{i}$ represent the NeRF loss function corresponding to the robot $i$, $\theta$ denotes the parameters of the 3D reconstruction model [25]. The relative poses are denoted by T, where the pose of the robot $i$ relative to the global coordinate system is represented by $T^{g}_{i}=[R|t]$ within the space of $SE(3)$; $g$ stands for global coordinate, which in this paper is the first robot coordinate ($T^{g}_{1}$ is the identity matrix). $N$ robots are connected with a communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consisting of vertices $\mathcal{V}=\{1,...,N\}$ and edges $\mathcal{E}$ which pairwise communication can occur. Here, the NeRF representation in [25] is used. Any other variants can also be used.
@@ III-B Relative Camera Pose Optimization
In multi-robot NeRF, robots do not send raw data over the wireless network. Therefore, using an SfM pipeline to calculate the relative poses for all robots is not easy. Here, the camera poses of each robot are expressed based on a local coordinate system, and gradually all poses transfer to the robot $R^{g}$ coordinate system (throughout this paper, the robot $R^{g}$ will be the arbitrary reference agent.). The optimization of the relative poses is performed jointly with the model parameters. Relative camera poses of multi-robot systems can be expressed as a camera-to-world transformation matrix (similar to each robot’s local camera pose) $T_{i}^{g}=[R|t]$in $SE(3)$, where $R\in SO(3)$ and $t\in\mathbb{R}^{3}$ show camera rotation and translation, respectively [20]. Optimizing the translation vector $t$ involves the designation of trainable parameters, given its definition in Euclidean space. For camera rotation, which is in $SO(3)$ space, the axis-angle representation is chosen: $\phi:=\alpha\omega,\phi\in\mathbb{R}^{3}$, where $\omega$ and $\alpha$ are a normalized rotation axis and a rotation angle, respectively. The rotation matrix $R$ can be expressed from Rodrigues’ formula:
$R=I+\frac{\sin{(\alpha)}}{\alpha}\phi^{\wedge}+\frac{1-\cos{(\alpha)}}{\alpha^{2}}(\phi^{\wedge})^{2}\ ,$ (2)
where $(.)^{\wedge}$ is the skew operator that converts a vector $\phi$ to a skew matrix. The relative camera poses for each robot $i$ are optimized by trainable parameters $\phi_{i}$ and $t_{i}$.
@@ III-C Distributed Formulation
In the problem presented in (1), each agent $i$ has its local cost function $L_{i}$, indicating that the optimization must be done in a distributed manner on each agent. However, $\theta$ is a global variable and is shared among all agents. This necessitates that the agents should achieve a consensus on the optimal value of $\theta$. In addition to the global variable $\theta$, agents have their local variables, such as $T_{i}^{g}$. To express that the optimization distributed,  (1) is rewritten as
$\theta,\textbf{T}$ $=\underset{\theta,\textbf{T}}{\operatorname{argmin}}\sum_{i\in\mathcal{V}}{L_{i}(\theta_{i},T^{g}_{i},I_{i},\pi_{i})}\leavevmode\nobreak\ ,$ (3)
@@
where each agent solves for its version of the global variable $\theta$, $\theta_{i}$, a relaxed local variable or primal variable. Assuming $\mathcal{N}_{i}$ is the set of neighbors of robot $i$, to achieve a consensus between agents $i$ and $\forall j\in\mathcal{N}_{i}$, an auxiliary variable $z_{ij}$ is introduced. This is also known as the ‘complicating variable’ which ensures that the agents achieve a consensus via ‘complicating constraints’: i.e. $\theta_{i}=z_{ij}$ and $\theta_{j}=z_{ij}$. The new problem in  (3) is then solved by introducing augmented Lagrangian and redefining the cost function to $\mathcal{L}_{i}$:
$\mathcal{L}_{i}(\theta_{i})=L_{i}(\theta_{i})+(\theta_{i}-z_{ij})^{\top}y_{i}+\frac{\rho}{2}\sum_{j\in\mathcal{N}_{i}}{\left\|\theta_{i}-z_{ij}\right\|_{2}^{2}}\ ,$ (4)
@@
There are various techniques to optimize  (4), such as auxiliary problem principle (APP) [48] and alternating direction method of multipliers (ADMM) [49]. ADMM has improved convergence properties [49]. We use a version of ADMM that is suitable for distributed systems, known as consensus ADMM (C-ADMM) [50] described next.

### v2

# Di-NeRF: Distributed NeRF for Collaborative Learning with Relative Pose Refinement

@@ Abstract
Collaborative mapping of unknown environments can be done faster and more robustly than a single robot. However, a collaborative approach requires a distributed paradigm to be scalable and deal with communication issues. This work presents a fully distributed algorithm enabling a group of robots to collectively optimize the parameters of a Neural Radiance Field (NeRF). The algorithm involves the communication of each robot’s trained NeRF parameters over a mesh network, where each robot trains its NeRF and has access to its own visual data only. Additionally, the relative poses of all robots are jointly optimized alongside the model parameters, enabling mapping with less accurate relative camera poses. We show that multi-robot systems can benefit from differentiable and robust 3D reconstruction optimized from multiple NeRFs. Experiments on real-world and synthetic data demonstrate the efficiency of the proposed algorithm. See the website of the project for videos of the experiments and supplementary material [^2] https://sites.google.com/view/di-nerf/home.
Index Terms:
Distributed Robot Systems; Mapping; Multi-Robot SLAM
@@ I Introduction
There is an increasing demand for robots to collaborate on complex tasks, such as mapping an unknown environment. Centralized approaches for the coordination of the robots or processing data face scalability issues, vulnerability to central node failures, and the risk of communication blackouts. A true collaborative robotic system needs to work in a distributed manner [1]. Further, building a high-quality representation in collaborative mapping is needed to make an informed decision. The representations also need to be compact for easy sharing. Neural Radiance Field (NeRF) [2] enables such a representation for a single robot, leveraging advancements in neural networks. Extending the capabilities of NeRF to facilitate multi-robot NeRF in a distributed manner emerges as a natural progression. This extension allows for a reliable and fast collaborative development of high-quality representations of unknown environments (e.g. in search & rescue, surveillance, and monitoring), without a central node. Performing distributed NeRF requires addressing several problems inherent to multi-robot systems [3]. Firstly, determining the relative poses of robots, preferably without requiring them to rendezvous. This is needed to enable the fusions of the local maps (i.e. maps of individual robots) into a global map. The relative poses can be determined either via line-of-sight rendezvous [4], which necessitates motion coordination, or by identifying overlaps within local maps [5], requiring additional processing. In large-scale applications, avoiding line-of-sight rendezvous is preferred to prevent further planning constraints. The second challenge is deciding what information to share among robots. While sharing unprocessed raw visual data is resource-intensive, depending on the architecture, sharing compact neural models is a more viable option. However, sharing maps introduces complexity to the subsequent challenge. The third challenge is merging representations, requiring the development of a method to unify local neural maps into a global map. Merging neural maps is more intricate than merging geometric representations.
@@
There are many methods for distributed/centralized geometric multi-robot mapping and localization [4, 6]. For learned methods, federated [7] and distributed learning [1] are common for overcoming the issues associated with centralized learning. In these methods, the training process is decentralized. Each robot performs its training, and the results are aggregated in a central node [7]. The need for a central node still constrains the scalability of such methods. A recent work using multiple agents for NeRF is Block-NeRF [8], composed of independent local maps, i.e. NeRF block. In Block-NeRF, the relative poses of the blocks are known to a degree and the task is done in a centralized manner. For large-scale aerial NeRF mapping, Mega-NeRF [9] partitions training images into NeRF submodules for parallel training. Another closely related work is DiNNO [10], which is a distributed algorithm enabling robots to collaboratively optimize deep neural networks over a mesh network without sharing raw data. In this paper, we present Di-NeRF (Distributed NeRF), an algorithm that builds on the Consensus Alternating Direction Method of Multipliers (C-ADMM) as a versatile distributed optimization method for multi-robot systems. In Di-NeRF, each robot starts building its NeRF by relying on its own data i.e. images and local camera poses from a Structure-from-Motion (SfM) algorithm. By integrating NeRFs from other robots, each robot will build a global NeRF model as if the robot has gathered all data and processed them. The advantages of Di-NeRF are that the robots do not know their exact initial poses, as illustrated graphically in Fig. 1. In Di-NeRF, robots alternate between local optimization of an objective function and communication of intermediate network weights over the wireless network. Di-NeRF can consider different communication graphs (e.g. fully connected, circular, and ring connectivity). The key contributions of Di-NeRF include i) developing fully distributed optimization for 3D reconstruction with RGB images as input, and a NeRF as the backend, enabling fusing NeRFs in training not the rendering process, and ii) optimization of the relative poses of the robots, removing the need to know the prior relative poses accurately. In this work, we make these assumptions: i) a reliable estimate of the agent’s trajectory in its local frame is available; ii) there is sufficient overlap between the agents’ viewpoints, allowing optimization of the relative poses.
@@ II Literature Review
Multi-robot localization and mapping algorithms utilize various representations such as sparse landmarks [11], dense geometric maps [12], object classes [13], and semantics [14]. With the advent of neural radiance field [2] and its variants [15, 16, 17, 18], there are efforts to use neural maps for collaborative localization/mapping [8, 19].
Emerging radiance fields, including non-neural representations such as Plenoxels [20], and neural representations, such as NeRF [2, 21] have revolutionized localization and mapping algorithms. Nice SLAM [22], iMAP [23], NeRF-SLAM [24], and [25] are using such representations. The early versions of NeRF representations [18], required pose information often generated via Structure-from-Motion (SfM) algorithms, e.g. COLMAP [26]. This limitation was later addressed in [15, 17, 27], by jointly optimizing camera poses and the scene. It was also shown that a camera view can be localized in a NeRF map, as shown in iNeRF [28].
Extending NeRF to multi-robot scenarios has also been proposed in Block-NeRF [8], NeRFuser [29], and NeRFusion [30]. These methods use blending techniques to render an image, often done centralized, performing inference on multiple NeRF models. Furthermore, it is assumed that there is prior knowledge about the relative pose of the robots. A similar research is federated learning [31, 19], where the training is done in a decentralized optimization, and each node learns a common NeRF in parallel. Then the weights are transferred to a server for aggregation. In DiNNO [10], a distributed algorithm for collaborative learning of a neural model is presented, demonstrating a range of experiments including multi-agent RL, MNIST classification, and 2D neural implicit mapping with known relative poses. The core of the algorithm is a distributed optimization method known as C-ADMM  [1].
@@ III Distributed Neural Radiance Field (Di-NeRF)
Here the problem is defined followed by the proposed method for relative pose refinement and distributed NeRF. The overall pipeline for Di-NeRF is shown in Fig. 2. Di-NeRF is a fully distributed algorithm for 3D reconstruction using RGB images. With Di-NeRF, each robot has its local coordinate and only shares the learned models to achieve a consensus while optimizing for the relative poses.
Figure 2: Three connected robots share their local NeRF to build a global NeRF. Each robot adjusts its weight compared to receiving weights using C-ADMM, and all robots except robot $R^{g}$ (Robot#1, the robot whose coordinate is considered as the global coordinate) optimize for the relative pose $T_{i}^{g}$ and depth distortion parameters.
@@ III-A Problem Statement
Assume $N$ robots connected to each other with a communication graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ consisting of vertices $\mathcal{V}=\{1,...,N\}$ and edges $\mathcal{E}$, on which pairwise communication can occur. Given a set of images $\boldsymbol{I}_{i}$ captured by robot $i$, $i=1,...,N$, from a scene with the associated local camera poses, $\boldsymbol{\pi}_{i}$, the goal of Di-NeRF is to build a scene representation, $\theta$, that enables all the robots to generate realistic images from novel viewpoints. Here, NeRF [18] is employed as the model for 3D reconstruction, $\theta$. Since the poses of each robot are defined in a local coordinate system, it is necessary to perform relative pose optimization to build a consistent model that accurately represents the entire scene. The problem is expressed as:
${\color[rgb]{0,0,0}\theta^{*}},{\color[rgb]{0,0,0}\boldsymbol{T}^{*}}=\underset{\theta,\mathbf{T}}{\operatorname{argmin}}\sum_{i\in\mathcal{V}}{L_{i}(\theta,T^{g}_{i}{\color[rgb]{0,0,0}\mid}\boldsymbol{I}_{i},\boldsymbol{\pi}_{i})},\leavevmode\nobreak\ \boldsymbol{T}=\{T^{g}_{1},..,T^{g}_{N}\},\leavevmode\nobreak\ (1)$
where $L_{i}$ represents the NeRF loss function corresponding to the robot $i$. The relative poses are denoted by $\boldsymbol{T}$, where the pose of the robot $i$ relative to the global coordinate system is represented by $T^{g}_{i}=[R^{g}|t^{g}]$ within the space of $SE(3)$. Superscript $g$ stands for the global coordinate. Without loss of generality, the first robot coordinate is selected as the global frame ($T^{g}_{1}$ is the identity matrix). If this robot fails, another robot is selected as the global frame. The model $\theta$ is shared among all robots and needs to be determined collectively, indicated by the summation in Eq. (III-A). Other variables with subscript $i$; i.e. $T^{g}_{i}$, $\boldsymbol{I}_{i}$, and $\boldsymbol{\pi}_{i}$; are local/internal to robot $i$. The model comprises fully connected neural network layers, described in Sec. IV-H. It learns to generate a color vector $\boldsymbol{c}=(r,g,b)$ and volume density $\sigma$ for each ray along a 3D location $\boldsymbol{x}=(x,y,z)$ and 2D viewing direction $\boldsymbol{d}$. Here, we use a pinhole camera model, assuming equal focal lengths for x and y directions. The camera ray that starts from camera origin $\boldsymbol{o}$ and passes through each pixel $\boldsymbol{p}=(u,v)$ is in the local coordinate of each robot and should be transferred to the global coordinate system using relative camera poses. In a local camera pose, $\pi_{i}=[R|t]$, the camera origin is $\boldsymbol{o}=t$ and the ray direction $\boldsymbol{d}$ is as follows [2, 18]:
$\boldsymbol{d}=R\begin{pmatrix}\frac{u-W/2}{f}&\frac{{\color[rgb]{0,0,0}-}v{\color[rgb]{0,0,0}+}H/2}{f}&-1\end{pmatrix}^{\top}\leavevmode\nobreak\ ,$ (2)
where $f$ is the focal length and $H$ and $W$ are the image height and width. By using the relative pose $T^{g}_{i}=[R^{g}|t^{g}]$, the camera pose in the global coordinate will be $\pi_{i}^{g}=T^{g}_{i}\pi_{i}=[R^{g}R|R^{g}t+t^{g}]$ which positions the camera origin at $R^{g}t+t^{g}$ and defines ray direction $R^{g}\boldsymbol{d}$. Since the relative poses, $\boldsymbol{T}^{g}_{i}$, are not known exactly, the model $\theta_{i}$ and relative poses must be optimized jointly. The optimization can be central, but a distributed approach is preferred for real-world use.
@@ III-B Relative Camera Pose Optimization
In multi-robot NeRF, robots do not send raw data over the wireless network. Therefore, using an SfM pipeline to calculate the relative poses for all robots is not easy. Here, the camera poses of each robot are expressed based on a local coordinate system, and gradually all poses transfer to the robot $R^{g}$ coordinate system (throughout this paper, the robot $R^{g}$ will be the arbitrary reference agent.). The optimization of the relative poses is performed jointly with the model parameters. Relative camera poses of multi-robot systems can be expressed as a camera-to-world transformation matrix (similar to each robot’s local camera pose) $T_{i}^{g}=[R^{g}|t^{g}]$in $SE(3)$, where $R^{g}\in SO(3)$ and $t^{g}\in\mathbb{R}^{3}$ show camera rotation and translation, respectively [15]. Optimizing the translation vector $t^{g}$ involves the designation of trainable parameters, given its definition in Euclidean space. For camera rotation, which is in $SO(3)$ space, the axis-angle representation is chosen: $\phi:=\alpha\omega,\phi\in\mathbb{R}^{3}$, where $\omega$ and $\alpha$ are a normalized rotation axis and a rotation angle, respectively. The rotation matrix $R^{g}$ can be expressed from Rodrigues’ formula:
$R^{g}=I+\frac{\sin{(\alpha)}}{\alpha}\phi^{\wedge}+\frac{1-\cos{(\alpha)}}{\alpha^{2}}(\phi^{\wedge})^{2}\ ,$ (3)
where $(.)^{\wedge}$ is the skew operator that converts a vector $\phi$ to a skew-symmetric matrix. The relative pose for each robot $i$ is optimized by trainable parameters $\phi_{i}$ and $t_{i}$.
Training NeRF and the relative camera poses jointly adds another degree of ambiguity to the problem, and causes the NeRF optimization to converge to local minima, especially for large relative poses. To avoid local minima, we use monocular depth estimation [17]. However, the depth maps suffer scale and shift distortions across frames, making mono-depth maps not multi-view consistent. To utilize mono-depth effectively, we explicitly optimize scale and shift parameters during NeRF training by penalizing the difference between rendered depth and mono-depth. For each image $I^{m}_{i}$ (out of $M_{i}$ images) of robot $i$, we generate mono-depth sequence $D^{m}_{i},m=0,...,M_{i}-1$ from input images with an off-the-shelf monocular depth network, i.e., DPT [32]. Two linear transformation parameters $\boldsymbol{\psi}_{i}^{m}=(\alpha^{m}_{i},\beta^{m}_{i}),m=0\ldots M_{i}-1$ for each mono-depth map as scale and shift factors can recover a multi-view consistent depth map ${D^{m}_{i}}^{*}=\alpha^{m}_{i}D^{m}_{i}+\beta^{m}_{i}$, for each image $m$ in each robot $i$. This joint optimization of distortion parameters and NeRF is achieved by adding a depth loss $L^{depth}_{i}=\sum_{m}||{D^{m}_{i}}^{*}-\hat{D}^{m}_{i}||$ to the RGB loss of NeRF [18], $L^{RGB}_{i}=\sum_{m}||{I^{m}_{i}}-\hat{I}^{m}_{i}||$. Here $\hat{D}^{m}_{i}$ and $\hat{I}^{m}_{i}$ are NeRF-rendered depth/RGB images, described in [17][18]. Thus, the local NeRF loss for each robot $i$ is defined as:
$L_{i}=L^{RGB}_{i}+\gamma L^{depth}_{i}\leavevmode\nobreak\ ,$ (4)
where $\gamma$ is a weighting factor for depth loss term. With depth information, Eq. (III-A) is rewritten as:
$\theta^{*},\boldsymbol{T}^{*},\boldsymbol{\psi}^{*}=\underset{\theta,\mathbf{T},\boldsymbol{\psi}}{\operatorname{argmin}}\sum_{i\in\mathcal{V}}{L_{i}(\theta,T^{g}_{i},\boldsymbol{\psi}_{i}\mid\boldsymbol{I}_{i},\boldsymbol{\pi}_{i})}\leavevmode\nobreak\ .$ (5)
$\boldsymbol{\psi_{i}}=\{{\psi}_{i}^{m}\}$ is the set of distortion parameters for robot $i$, and $\boldsymbol{\psi}=\{\boldsymbol{\psi}_{i}\}$ is the set of all distortion parameters of all robots.
@@ III-C Distributed Formulation
In the problem presented in Eq. (5), each agent $i$ has its local cost function $L_{i}$, indicating that the optimization must be done in a distributed manner on each agent. However, $\theta$ is a global variable and is shared among all agents. This necessitates that the agents should achieve a consensus on the optimal value of $\theta$. In addition to the global variable $\theta$, agents have their local variables, such as $T_{i}^{g}$. To express that the optimization is distributed, Eq. (5) is rewritten as
${\color[rgb]{0,0,0}\theta^{*},\boldsymbol{T}^{*},\boldsymbol{\psi}^{*}}$ $=\underset{\theta,\boldsymbol{T},{\color[rgb]{0,0,0}\boldsymbol{\psi}}}{\operatorname{argmin}}\sum_{i\in\mathcal{V}}L_{i}(\underbrace{\theta_{i},T_{i}^{g},{\color[rgb]{0,0,0}\boldsymbol{\psi}_{i}}}_{{{\color[rgb]{0,0,0}\Gamma_{i}}}}{\color[rgb]{0,0,0}\mid}\boldsymbol{I}_{i},\boldsymbol{\pi}_{i})\leavevmode\nobreak\ .$ (6)
@@
Each agent solves for its version of the global variable $\theta$, $\theta_{i}$, a relaxed local variable or primal variable. Assuming $\mathcal{N}_{i}$ is the set of neighbors of robot $i$, to achieve a consensus between agents $i$ and $\forall j\in\mathcal{N}_{i}$, an auxiliary variable $z_{ij}$ is introduced. This is also known as the ‘complicating variable’ which ensures that the agents achieve a consensus via ‘complicating constraints’: i.e. $\theta_{i}=z_{ij}$ and $\theta_{j}=z_{ij}$. By defining ${\Gamma}_{i}=\{\theta_{i},T_{i}^{g},\boldsymbol{\psi}_{i}\}$ for notation brevity, Eq. (6) is then solved by introducing augmented Lagrangian and redefining the cost function to $\mathcal{L}_{i}$ – for brevity, $\boldsymbol{I}_{i}$ and $\boldsymbol{\pi}_{i}$ were omitted:
[…]
@@
There are various techniques to optimize Eq. (7), such as auxiliary problem principle (APP) [33] and alternating direction method of multipliers (ADMM) [34]. ADMM has improved convergence properties [34]. We use a version of ADMM that is suitable for distributed systems, known as consensus ADMM (C-ADMM) [35] described next.